how many kg of dinitrogen monoxide are present in 6.17 L of gas at STP? (scientific notation)
1
Expert's answer
2020-12-21T03:52:09-0500

m(N2O)= (V/Vm) * Mr= 6.17/22.4*44= 12.12 g


Answer: 12.12 g